A 900N crate slides 15m down a ramp that makes an angle of 35 degrees with the horizontal.

If the crate slides at a constant speed, how much thermal energy is created?
Answer #1

that means all graviational energy goes to thermal energy


so thermal = m g h = 900*12*sin(40 degrees)=6942 J
Answer #2

delta E = mgl sintheta

=wlsintheta

= 900*12*sin40 = 6942 J
Answer #3

as it slides at constant speed, so frictional force must be equal to mgsin35

so, F= 900*sin40=578.51 N

so thermal energy created =F*12 =6942.11 J = 6.94 kJ
Answer #4

E = 900*15*sin(35)

E = 7743 J